Homework Problem 43 kN/m А с The compound beam is pin supported at B and supported by rockers at A and C. There is a hinge (pin) at D. Determin
Answer #1

ACCORDING TO THE DATA AND DIAGRAM ARE GIVEN IN THE QUESTION,

Free body diagram : 3(11) 3(3) Dy V B AC Bx 1.5 1.5m 1:5 m 5.5m 4m N. Du ву. NA The compound beam is dismembered into members+ Fe = 0; Dox=0 Now, Member ABD, (+ EMBE B = 0; NA (7) + Dy (4) = 3(11) (1-5) 3(11)(1.5) – 4(4.5) Na (7) NA = 4.5 KN Ans By (